The length of a rectangle is 2.5 cm longer than twice the width. The perimeter is 40 cm. What are the dimensions of the rectangle
Answer by mananth(16946)

width = w
length = 2w+2.5
..
2*(L+W)= Perimeter
2*(w+2w+2.5)=40
2*(3W+2.5)=40
/2
3W+2.5 = 20
-2.5
3W= 20-2.5
3W= 17.5
W=5.83 cm the width
...
L= 2W+2.5
L=2*5.83+2.5
L=11.66+2.5
L=14.16 cm the length
...
m.ananth@hotmail.ca